OpenStudy (anonymous):

Can someone help me finish this equation? I'm using the quadratic equation. g(x) = x^2 + 6x + 1 changed it to x2 +6x+1=0 So then I plugged it in; x = [-6 (plus minus sign) sqrt 6^2 - 4(1)(1) ] /2 x = [ -6 (plus minus sign) sqrt 36 - 4] /2 x = [-6 (plus minus sign) sqrt (32)]/2 I know the sqrt of 32 is 4sqrt 2 but I don't know how to include it in there. Please help, and you can include any mistakes I might have made too :D

OpenStudy (tkhunny):

?? \(\dfrac{-6 \pm 4\sqrt{2}}{2} = \dfrac{2(-3\pm2\sqrt{2})}{2} = \dfrac{2}{2}(-3\pm2\sqrt{2}) = -3\pm2\sqrt{2}\)

OpenStudy (anonymous):

Oh that was simpler than I thought o-o thank you so much!! I over think things so it all was jumbled up in my head xD